# Layer by Layer: Uncovering Hidden Representations in Language Models

### v1

@@ Abstract
From extracting features to generating text, the outputs of large language models (LLMs) typically rely on their final layers, following the conventional wisdom that earlier layers capture only low-level cues. However, our analysis shows that intermediate layers can encode even richer representations, often improving performance on a wide range of downstream tasks. To explain and quantify these hidden-layer properties, we propose a unified framework of representation quality metrics based on information theory, geometry, and invariance to input perturbations. Our framework highlights how each model layer balances information compression and signal preservation, revealing why mid-depth embeddings can exceed the last layer’s performance. Through extensive experiments on 32 text-embedding tasks and comparisons across model architectures (transformers, state-space models) and domains (language, vision), we demonstrate that intermediate layers consistently provide stronger features. These findings challenge the standard focus on final-layer embeddings and open new directions for model analysis and optimization, including strategic use of mid-layer representations for more robust and accurate AI systems.
@@ 1 Introduction
Large Language Models (LLMs) have driven remarkable progress in natural language processing (NLP), achieving state-of-the-art results on many tasks (Brown et al., 2020; Devlin, 2018; Li et al., 2022). At the heart of most applications lies a common assumption: final-layer representations are the most useful for downstream tasks. Yet a fundamental question remains: does the final layer always yield the best representation?
In this paper, we conduct a layer-wise analysis of LLMs across diverse architectures—including Transformers (Vaswani et al., 2017), state-space models (SSMs) (Gu & Dao, 2024), and encoder-based models like BERT (Devlin, 2018)—spanning parameter scales from tens of millions to billions. Through systematic evaluation on 32 embedding tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), we find that intermediate layers often surpass the final layer by up to 16% in downstream accuracy. Figure 1 illustrates this phenomenon, where mid-depth layers provide particularly strong representations while the very last layer can become overly specialized to the pretraining objective.
A unified framework. To understand intermediate layers’ effectiveness, we integrate three complementary perspectives (Section 3):
@@
  • Intermediate layers consistently outperform final layers. This pattern is evident not only in Transformers but also in SSMs, suggesting a broad, architecture-agnostic effect.
@@
Overall, our results challenge the default reliance on final-layer embeddings and highlight intermediate layers as potentially underutilized sources of meaningful features. In the rest of this paper, we detail our unified framework (Section 3), present an extensive set of experiments in both language and vision (Section 4, 6), and conclude with a discussion of implications for model design, training practices, and future directions.
@@ 2 Related Work
A long line of research has aimed to understand how deep neural networks encode and organize information. Early studies employed linear probes to interpret intermediate layers (Alain & Bengio, 2017), while subsequent efforts introduced more sophisticated techniques such as SVCCA (Raghu et al., 2017) to compare learned features across architectures and training regimes. Although these approaches shed light on representation dynamics, most focus on vision backbones or relatively shallow models. In contrast, our work extends layer-wise analysis to large-scale language models, highlighting specific behaviors of intermediate layers in autoregressive Transformers, state-space models (SSMs), and beyond.
@@
Transformer-based LLMs have sparked significant interest in which layers capture linguistic properties such as syntax and semantics (Liu et al., 2019; Tenney et al., 2019; Voita et al., 2019). More recent work (Jin et al., 2024; Gurnee & Tegmark, 2023; Fan et al., 2024) has shown that mid-depth layers sometimes hold surprisingly robust features, challenging the typical focus on final layers. Our contribution unifies and expands these observations through a large-scale, theoretical–empirical framework that quantifies the quality of every layer’s representation via information theory, geometry, and invariance metrics.
Architectural Comparisons.
Transformers remain the dominant architecture for NLP (Vaswani et al., 2017), but they come in multiple variants. Encoder-only models (e.g., BERT (Devlin, 2018)) typically use bidirectional attention and masked-language objectives, while decoder-only architectures (e.g., GPT (Brown et al., 2020)) follow an autoregressive paradigm. Meanwhile, newer state-space models (SSMs) such as Mamba (Gu & Dao, 2024) use recurrent-style dynamics for efficient long-sequence processing. Although these designs differ significantly in attention mechanisms and sequence modeling strategies, there has been little direct comparison of hidden-layer representations across them. In our work, we analyze Transformers (both encoder- and decoder-only) and SSMs under a common set of metrics, highlighting contrasts in how intermediate layers compress or preserve information and showing that intermediate-layer representations can excel across multiple architectures.
@@
  • Information-theoretic measures capture how much a model’s internal representations compress or preserve relevant information. For example, the Information Bottleneck (Shwartz-Ziv & Tishby, 2019; Shwartz-Ziv, 2022) analyzes whether intermediate layers discard noise while retaining essential features.
  • Geometric measures focus on the structure of embeddings in high-dimensional space. Classical approaches include analyzing singular values or effective rank of the representation matrix (Garrido et al., 2023), while more recent work explores curvature (Hosseini & Fedorenko, 2023) to quantify how smoothly tokens are mapped across consecutive positions or time steps.
  • Task-based or invariance metrics evaluate how well representations support downstream goals. For instance, augmentations-based approaches such as InfoNCE (Oord et al., 2018) and LiDAR (Thilak et al., 2024) estimate invariance to perturbations, while methods like NESum or Self-Cluster (Agrawal et al., 2022) link closely to entropy. In computer vision, these scores often correlate strongly with downstream accuracy, highlighting how robust the embeddings are.
Although these categories may appear distinct, we will show (Section 3) that many can be unified under a single lens. This unification illuminates why certain intermediate layers balance compression, geometry, and invariance so effectively, leading to better representations for downstream tasks.
Compression and Generalization.
Multiple lines of research connect compression and generalization performance (Deletang et al., 2024). For instance, Bordes et al. (2023) demonstrated that discarding certain layers in self-supervised encoders can even improve downstream accuracy, while Park et al. (2024a) found that LLM embeddings often lie in low-dimensional manifolds. Our empirical study reinforces these ideas by demonstrating that many networks—especially autoregressive Transformers—naturally develop a mid-layer bottleneck that appears crucial for balancing “signal” versus “noise.” We show how intermediate layers can achieve optimal trade-offs between preserving task-relevant information and discarding superfluous detail.
@@ 3.1 Notation and Motivation
    How robust are they to small perturbations or augmentations?
@@
The answers can illuminate which layers strike the right balance between preserving relevant features and discarding noise.
@@ 3.2 Matrix-Based Entropy: A Common Theoretical Thread
Implication: bridging geometry, invariance, and local vs. global features.
@@ 3.3.1 Information-Theoretic Metrics
Effective Rank  (Roy & Vetterli, 2007a)
can be shown to be a lower bound to $\exp(S_{1}(\mathbf{Z}))$, highlighting how dimensionality effectively shrinks if the representation is strongly compressed. We prove this connection later in Theorem 1. This has implications for popular representation evaluation metrics such as RankMe (Garrido et al., 2023) and LiDAR (Thilak et al., 2024), which are both inspired by Effective Rank.
@@ 3.3.3 Augmentation Invariance Metrics
Lastly, we assess how stable the model’s representations are to small perturbations of the same input (e.g., random character swaps, keyboard-level changes; see Appendix). Suppose $p_{i}$ is augmented into $p_{i}^{(a)}$ and $p_{i}^{(b)}$. After embedding these, we compare the row vectors in $\mathbf{Z}_{1},\mathbf{Z}_{2}\in\mathbb{R}^{N\times D}$ under different scoring criteria:
@@ 3.4 Core Theoretical Results
Here, we summarize key statements that justify why these metrics meaningfully measure representation quality. We refer to the appendix F for details and proofs. Beyond serving as a unifying view, matrix-based entropy also connects to foundational concepts like majorization, Schur concavity, and mutual information. Furthermore, we can directly relate the eigenvalue entropy to the matrix entropy, most naturally via the Effective Rank (Roy & Vetterli, 2007b). The following theorem makes this connection explicit.
@@
    If prompt entropy remains near its maximum for all prompts, then the dataset entropy $S_{2}\!\bigl{(}\overline{\mathbf{Z}}\,\overline{\mathbf{Z}}^{\top}\bigr{)}$ grows on the order of $\log\!\bigl{(}\tfrac{D^{2}}{N}\bigr{)}.$
@@
    If prompt entropy instead stays near its minimum for all prompts, then dataset entropy grows more slowly, on the order of $\log\!\bigl{(}\tfrac{D^{2}}{N^{3}}\bigr{)}.$
@@ 4 Empirical Results
In this section, we empirically validate our theoretical framework through extensive experiments across architectures, scales, and training regimes. Our investigation centers on three key questions:
@@ 4.1.1 Experimental Setup
We evaluate three distinct architectural families: Pythia and Llama3 (decoder-only transformer) (Biderman et al., 2023; Dubey et al., 2024), Mamba (state space model) (Gu & Dao, 2024), BERT (encoder-only transformer) (Devlin, 2018) and LLM2Vec models (bidirectional attention) (BehnamGhader et al., 2024).
@@
We test each layer’s embeddings on 32 tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), spanning classification, clustering, and reranking. This comprehensive evaluation provides insight into how different layers capture task-relevant features. For a full list of the tasks, refer to the Appendix.
@@ 4.1.2 Intermediate Layers Often Outperform Final Layers
A key question is whether final-layer embeddings are indeed optimal for downstream tasks. In Figure 1, we compare average performance on MTEB tasks across all layers of the three models.
@@ 4.1.3 Layer-Wise Metrics Correlate with Downstream Performance
To validate our framework’s relevance, we analyze how each metric (entropy, InfoNCE, etc.) correlates with downstream performance. Figure 3 and Figure 8 show distance correlations between metrics and task scores for Pythia-410M. We make several key observations:
  • All metrics show strong relationships with performance
  • DiME, curvature, and InfoNCE exhibit particularly strong correlations
  • Associations remain robust across different correlation measures (Spearman, Kendall)
These relationships suggest that our metrics effectively capture what makes intermediate representations powerful for downstream tasks.
@@ 4.2 Architectural and Scale Differences
Figure 2 shows how prompt entropy, curvature, and augmentation metrics evolve across each model’s layers. BERT, which encodes the entire input bidirectionally, generally maintains high entropy across layers, suggesting minimal compression: the model can see all tokens at once and need not discard as much information. By contrast, the decoder-only Pythia exhibits a strong mid-layer entropy dip, reflecting its autoregressive objective’s tendency to filter or prune non-local details in the middle of the network. As a result, Pythia’s “sweet spot” for downstream tasks often lies around mid-depth, where it balances essential context and compression. Mamba, meanwhile, processes sequences through a state-space approach that yields flatter, more uniform curves across depth: it neither retains as much information as BERT nor compresses as aggressively as Pythia’s mid-layers.
@@
In Figure 12, we analyze Pythia models ranging from 14M to 1B parameters. Larger Pythia models display more pronounced intermediate compression (entropy dips), indicating a heightened ability to distill relevant features. We also observe smoother token trajectories (lower curvature) and stronger invariance (e.g., higher LiDAR), consistent with findings that bigger models more effectively filter noise and capture long-range dependencies. Notably, these trends further reinforce why performance often peaks in the middle of the network: larger models gain more capacity to compress intermediate representations, yet still preserve crucial semantic details.
@@
While our experiments treat each Transformer layer as a single unit, Transformer blocks comprise multiple sub-layers (pre-attention normalization, self-attention, residuals, MLPs). By measuring entropy after each sub-layer, we find in Figure 15 that residual connections drive the mid-network compression observed in Section 4.2. Specifically:
  • Sub-layers before residuals (e.g. pre-attention, raw attention, or MLP pre-residual outputs) often show only mild compression; their representations still carry much of the original variability
  • Residual sub-layers exhibit a marked entropy drop, reflecting significant information filtering
The strong “valley” in entropy at intermediate layers is thus tied to how residual paths merge computed signals with the existing hidden state. This aligns with prior work indicating that residuals act as a regularizer or "noise filter" (Marion et al., 2024), smoothing out spurious components in hidden representations.
@@ 4.3 Impact of Training Progression
We measure Pythia’s metrics at multiple checkpoints to understand how layer-wise representations evolve throughout training (Figure 4). Two main observations emerge:
@@
In contrast to the intermediate layers, the earliest layers change very little after the initial phase of training. This observation aligns with the “detokenization” hypothesis (Lad et al., 2024), which posits that early layers mainly convert raw tokens into a basic embedding space and then remain relatively fixed. As a result, the most substantial improvements in representation quality—such as enhanced compression—are driven primarily by the intermediate layers, reinforcing their importance for learning robust, high-level features.

### v2

@@ Abstract
From extracting features to generating text, the outputs of large language models (LLMs) typically rely on the final layers, following the conventional wisdom that earlier layers capture only low-level cues. However, our analysis shows that intermediate layers can encode even richer representations, often improving performance on a range of downstream tasks. To explain and quantify these hidden-layer properties, we propose a unified framework of representation quality metrics based on information theory, geometry, and invariance to input perturbations. Our framework highlights how each layer balances information compression and signal preservation, revealing why mid-depth embeddings can exceed the last layer’s performance. Through extensive experiments on 32 text-embedding tasks across various architectures (transformers, state-space models) and domains (language, vision), we demonstrate that intermediate layers consistently provide stronger features, challenging the standard view on final-layer embeddings and opening new directions on using mid-layer representations for more robust and accurate representations.
@@ 1 Introduction
Large Language Models (LLMs) have driven remarkable progress in natural language processing (NLP), achieving state-of-the-art results on many tasks (Brown et al., 2020; Devlin et al., 2019; Li et al., 2022). At the heart of most applications lies a common assumption: final-layer representations are the most useful for downstream tasks. Yet a fundamental question remains: does the final layer always yield the best representation?
In this paper, we conduct a layer-wise analysis of LLMs across diverse architectures—including transformer-based ones (Vaswani et al., 2017), state-space models (SSMs) (Gu & Dao, 2024), and encoder-based models like BERT (Devlin et al., 2019)—spanning parameter scales from tens of millions to billions. Through systematic evaluation on 32 embedding tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), we find that intermediate layers often surpass the final layer by up to 16% in downstream accuracy. Figure 1 illustrates this phenomenon, where mid-depth layers provide particularly strong representations while the very last layer can become overly specialized to the pretraining objective.
A unified framework. To better understand intermediate layers’ effectiveness, we combine three complementary perspectives (Section 3):
@@
  • Intermediate layers consistently outperform final layers. This pattern is evident in both transformers and SSMs, suggesting a broad architecture-agnostic effect.
@@
Overall, our results challenge the default reliance on final-layer embeddings and highlight intermediate layers as potentially underutilized sources of meaningful features. In this paper, we detail our unified framework (Section 3), present extensive experiments in both language and vision (Section 4, 5, 6), and conclude with a discussion of our findings, their implications, and future directions.  [^1] We make our code available at https://github.com/OFSkean/information_flow
@@ 2 Related Work
A long line of research has aimed to understand how deep neural networks encode and organize information. Early studies employed linear probes for intermediate layers (Alain & Bengio, 2017), while subsequent efforts introduced more sophisticated techniques such as SVCCA (Raghu et al., 2017) to compare learned features across architectures and training regimes. Although these approaches have shed light on representation dynamics, most focus on vision models or shallow networks. In contrast, our work contributes to a growing body of literature extending layer-wise analysis to large-scale language models, emphasizing specific behaviors of intermediate layers across diverse architectures. Complementing our empirical findings, Saponati et al. (2025) present a theoretical analysis of how different pretext tasks, such as next-token prediction and masked language modeling, influence the structure of learned representations.
@@
Recent work has increasingly focused on identifying which transformer layers encode different types of information. For example, linguistic features such as part-of-speech tags or semantic roles are best encoded by the middle layers of a BERT (Liu et al., 2019; Tenney et al., 2019; Voita et al., 2019). More recent work has shown that mid-depth layers sometimes hold surprisingly robust features, challenging the usual emphasis on final layer representations (Jin et al., 2024; Gurnee & Tegmark, 2023; Fan et al., 2024). A related line of work investigates the attention sink phenomenon (Xiao et al., 2024; Brunner et al., 2020; Gu et al., 2025), in which attention disproportionately concentrates on a single token. Notably, intermediate decoder layers have been shown to not exhibit these extreme attention sinks (Barbero et al., 2025), suggesting they engage in more distributed and meaningful information processing than the shallow or deep layers.
Compression and Generalization.
Multiple lines of research link compression and generalization performance (Deletang et al., 2024). For instance, Bordes et al. (2023) demonstrated that discarding certain layers in self-supervised encoders can even improve downstream accuracy, while Park et al. (2024a) found that LLM embeddings often lie in low-dimensional manifolds. Our empirical study reinforces these ideas by demonstrating that many networks—especially autoregressive transformers—naturally develop a mid-layer bottleneck that appears crucial for balancing “signal” versus “noise.” We show how intermediate layers can achieve optimal trade-offs between preserving task-relevant information and discarding superfluous detail.
@@
  • Information-theoretic measures capture how much a model’s internal representations compress or preserve relevant information. For example, the Information Bottleneck (Shwartz-Ziv & Tishby, 2019; Shwartz-Ziv, 2022) analyzes whether intermediate layers discard noise while retaining essential features. Intrinsic dimensionality, which describes the minimum number of features to represent data, has also been used to analyze intermediate layers in LLMs (Cheng et al., 2025; Valeriani et al., 2023; Razzhigaev et al., 2024). This line of work has shown semantic abstractions useful for downstream tasks are better encoded in middle layers than last layers in large transformer models. While we do not study intrinsic dimensionality in our subsequent analysis, it would make a promising direction for future work.
  • Geometric measures focus on the structure of embeddings in high-dimensional space. Classical approaches include analyzing singular values and effective rank of the representation matrix (Garrido et al., 2023). The anisotropy metric of Razzhigaev et al. (2024) has been used to study compression in intermediate model layers and we compare our results with their findings in Section 4.2. Anisotropy fits in well with our proposed framework, though we leave a formal integration to future work. Recent work explores curvature (Hosseini & Fedorenko, 2023) to quantify how smoothly tokens are mapped across consecutive positions or time steps.
  • Task-based or invariance metrics evaluate how well representations support downstream goals. For instance, augmentations-based approaches such as InfoNCE (Oord et al., 2018) and LiDAR (Thilak et al., 2024) estimate invariance to perturbations, while methods like NESum or Self-Cluster (Agrawal et al., 2022) link closely to entropy. In computer vision, these scores often correlate strongly with downstream accuracy, highlightFing how robust the embeddings are.
Although these representation quality metric categories may appear distinct, we show (Section 3) that many can be unified under a single lens. This unification illuminates why certain intermediate layers balance compression, geometry, and invariance so effectively, leading to better representations for downstream tasks.
@@ 3.1 Notation and Motivation
    How robust are they to perturbations or augmentations?
@@
Answers to these questions can illuminate which layers strike the right balance between preserving relevant features and discarding noise.
@@ 3.2 Matrix-Based Entropy: A Common Theoretical Thread
Bridging geometry, invariance, and feature locality.
@@ 3.3.1 Information-Theoretic Metrics
Effective Rank.
(Roy & Vetterli, 2007) can be shown to be a lower bound to $\exp(S_{1}(\mathbf{Z}))$, highlighting how dimensionality effectively shrinks if the representation is strongly compressed. We prove this connection in Theorem 1. This has implications for popular representation evaluation metrics such as RankMe (Garrido et al., 2023) and LiDAR (Thilak et al., 2024), which are both inspired by Effective Rank.
@@ 3.3.3 Augmentation Invariance Metrics
Lastly, we assess how stable the model’s representations are to small perturbations of the same input (e.g., random character swaps, keyboard-level changes; see Appendix). Suppose a prompt $p_{i}$ is augmented into $p_{i}^{(a)}$ and $p_{i}^{(b)}$. After embedding these, we compare the row vectors in $\mathbf{Z}_{1},\mathbf{Z}_{2}\in\mathbb{R}^{N\times D}$ under different scoring criteria:
@@ 3.4 Core Theoretical Results
Here, we summarize key statements that justify why these metrics meaningfully measure representation quality. We refer to the appendix G for details and proofs. Beyond serving as a unifying view, matrix-based entropy also connects to foundational concepts like majorization, Schur concavity, and mutual information. Furthermore, we can directly relate the eigenvalue entropy to the matrix entropy, most naturally via the Effective Rank (Roy & Vetterli, 2007). The following theorem makes this connection explicit.
@@
    If prompt entropy remains near its maximum for all prompts, then the dataset entropy $S_{2}\!\bigl{(}\overline{\mathbf{Z}}\,\overline{\mathbf{Z}}^{\top}\bigr{)}$ grows on the order of $\log\!\bigl{(}\tfrac{L^{2}}{N}\bigr{)}.$
@@
    If prompt entropy instead stays near its minimum for all prompts, then dataset entropy grows more slowly, on the order of $\log\!\bigl{(}\tfrac{L^{2}}{N^{3}}\bigr{)}.$
@@ 4 Empirical Results
In this section, we empirically test our theoretical framework through extensive experiments across architectures, scales, and training regimes. We focus on three key questions:
@@ 4.1.1 Experimental Setup
We evaluate three distinct architectural families: Pythia and Llama3 (decoder-only transformers) (Biderman et al., 2023; Dubey et al., 2024), Mamba (state space model) (Gu & Dao, 2024), BERT (encoder-only transformer) (Devlin et al., 2019) and LLM2Vec models (bidirectional attention) (Behnam Ghader et al., 2024).
@@
We test each layer’s embeddings on 32 tasks from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), spanning classification, clustering, and reranking dor a comprehensive evaluation across various tasks. We refer to the Appendix for details.
@@ 4.1.2 Intermediate Layers Often Outperform Final Layers
Are final-layer embeddings indeed optimal for downstream tasks? In Figure 1, we compare average performance on MTEB tasks across all layers of the three models.
@@ 4.1.3 Layer-Wise Metrics Correlate with Downstream Performance
To validate our framework, we analyze how each evaluation metric correlates with downstream performance. Figures 3 and 8 show distance correlations between metrics and task scores for Pythia-410M. We find that all metrics exhibit strong relationships with downstream performance. Among them, curvature, DiME, and InfoNCE stand out with particularly high correlations. These associations remain robust across different correlation measures, including Spearman and Kendall, reinforcing the reliability of our findings.
Our results suggest that our metrics capture some aspects of intermediate representations that contribute to downstream utility. In Appendix E, we leverage these strong correlations to select high-performing layers in an unsupervised manner, following (Agrawal et al., 2022; Garrido et al., 2023; Thilak et al., 2024). In short, we can identify an intermediate layer that surpasses the final layer in downstream performance—without using any task-specific labels. For instance, using DiME-based layer selection leads to a 3% average improvement in MTEB scores for the Pythia-410M model.
@@ 4.2 Architectural and Scale Differences
Figure 2 shows how prompt entropy, curvature, and augmentation metrics evolve across each model’s layers. BERT, which encodes the entire input bidirectionally, generally maintains high entropy across layers, suggesting minimal compression: the model can see all tokens at once and need not discard as much information. By contrast, the decoder-only Pythia exhibits a strong mid-layer entropy dip, reflecting its autoregressive objective’s tendency to filter or prune non-local details in the middle of the network. As a result, Pythia’s “sweet spot” for downstream tasks often lies around mid-depth, where it balances essential context and compression. Mamba, meanwhile, processes sequences through a state-space approach that yields flatter, more uniform curves across depth: it neither retains as much information as BERT nor compresses as aggressively as Pythia’s mid-layers. These conclusions align with  Razzhigaev et al. (2024) which showed a flat layer-wise anisotropy for encoder models and a spike in intermediate layer anisotropy for decoder models.
@@
In Figure 12, we analyze Pythia models ranging from 14M to 1B parameters. Larger models display more pronounced intermediate compression (entropy dips), indicating a heightened ability to distill relevant features. We also observe smoother token trajectories (lower curvature) and stronger invariance (higher LiDAR), consistent with findings that bigger models more effectively filter noise and capture long-range dependencies. These trends reinforce why performance peaks in the middle of the network: larger models hold more capacity to compress intermediate representations, yet still preserve crucial semantic details.
@@
While our experiments treat each transformer layer as a single unit, transformer blocks are composed of multiple sub-layers (pre-attention normalization, self-attention, residuals, MLPs). By measuring entropy after each sub-layer, we find in Figure 15 that residual connections drive the mid-network compression observed in Section 4.2. Specifically:
  • Sub-layers before residuals (e.g. pre-attention, attention scores, or MLP pre-residual outputs) often show only mild compression.
  • Residual sub-layers exhibit a pronounced drop in entropy, indicating a significant filtering of information. A concurrent study (Csordás et al., 2025) observed a decrease in the residual stream norm in the second half of decoder models, reinforcing our findings.
The strong entropy “valley” at intermediate layers is tied to how residual paths merge new signals with the existing hidden state. This aligns with prior work indicating that residuals act as a regularizer (Marion et al., 2024), smoothing out spurious components in hidden representations.
@@ 4.3 Impact of Training Progression
We measure Pythia’s metrics at multiple checkpoints to understand how layer-wise representations evolve throughout training (Figures 4 and  11). Two main observations emerge:
@@
In contrast to intermediate layers, the earliest layers change very little after the initial phase of training. This observation aligns with the “detokenization” hypothesis of Lad et al. (2024), which posits that the main functional role of early layers is to convert raw tokens into a basic embedding space. This idea is closely related to the “shared task” layers of  Zhao et al. (2024), introduced in the context of instruction tuning on diverse tasks. In particular, they show that the first nine layers of LlaMA 2 7B (Touvron et al., 2023) perform general task-agnostic operations. As a result, the most substantial changes to representations, such as enhanced compression, are driven primarily by the intermediate layers, reinforcing their importance for learning robust, high-level features.